OpenStudy (agentjamesbond007):

[Related Rates] If A is the area of a circle with radius r and the circle expands as time passes, find dA/dt in terms of dr/dt

OpenStudy (anonymous):

like the last one \[A=\pi r^2\] \[\frac{dA}{dt}=\frac{dA}{dr}\frac{dr}{dt}=2\pi r\frac{dr}{dt}\]